OpenStudy (anonymous):

SOlve using the quadratic formula 4x²+7x+1=0

OpenStudy (ddcamp):

Given an equation in the form: \[ax^2 + bx + c = 0\] The quadratic formula is: \[x = \frac{ -b ± \sqrt{b²-4ac} }{ 2a }\]

OpenStudy (anonymous):

we can find the discriminant d=\[b ^{2}-4ac\] we get d =33 now as d>0 we have two real values of x using the formula +-x=\[-B \pm \sqrt{d}/2a\]